# Identifying Network Structure of Linear Dynamical Systems: Observability and Edge Misclassification

### v1

@@ III-A Equivalence of the $\ell_{1}$ and $\ell_{2}$ Programs
Relaxing the program given in (13) to minimizing the $\ell_{2}$ norm, we arrive at the following optimization program
@@
As this problem is convex, to find the optimal solution it suffices to take the gradient of the objective yielding
@@
We note that this is a linear equation when in vectorized form, namely $(I_{n}\otimes\Phi^{\top})(a+\Gamma v^{*})=0$. Given the solution to this matrix equation, $V^{*}$, we now provide a condition under which this solution is also a solution to (13).
@@
The solution, $V^{*}$, to (17) is a solution to the original program (13) if
@@
Let $v^{*}=\text{vec}(V^{*})$ where $V^{*}$ is a solution to (17). Moreover, assume that $\Phi^{\top}(Z\odot\text{sgn}(A+Z\odot(\Phi V^{*})))=0$. Recall that the programs (13) and (16) are equivalent, hence we show that for an appropriate choice $(v^{*},t^{*},\lambda^{*})$, where $\lambda^{*}=\begin{bmatrix}{\lambda_{1}^{*}}^{\top}&{\lambda_{2}^{*}}^{\top}\end{bmatrix}^{\top}$ is the dual variable associated with the constraint in (16), the Karush-Kuhn-Tucker (KKT) conditions for (16) are satisfied. We select $t^{*}=|a+\Gamma v^{*}|$, and $\lambda_{1,i}^{*}=1$ and $\lambda_{2,i}^{*}=0$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=1$, $\lambda_{1,i}^{*}=0$ and $\lambda_{2,i}^{*}=1$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=-1$, and lastly $\lambda_{1,i}^{*}=1/2=\lambda_{2,i}^{*}$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=0$. With these choices $\lambda_{1}^{*}-\lambda_{2}^{*}=\text{sgn}(a+\Gamma v^{*})$. The Lagrangian of (16) is
@@
$\begin{bmatrix}0\\ \mathbbm{1}\end{bmatrix}+\begin{bmatrix}\Gamma^{\top}&-\Gamma^{\top}\\ -I_{n^{2}}&-I_{n^{2}}\end{bmatrix}\lambda^{*}=0$ (21a)
@@
With $t^{*}$ as chosen, the feasibility condition (21b) is satisfied. Moreover, (21c) is satisfied by our choice of $\lambda^{*}$. It remains to check (21a) and (21d). Notice, from the stationarity of the Lagrangian (21a) we have that $\lambda_{1}^{*}+\lambda_{2}^{*}=\mathbbm{1}$ which is also satisfied from our choice of $\lambda^{*}$. Rewriting (21d) as
@@
Since $\lambda_{1}^{*}-\lambda_{2}^{*}=\text{sgn}(a+\Gamma v^{*})$ these equations are satisfied, and thus (21d) is satisfied. Lastly, we must consider the first n(nr)n(n-r) rows of the stationarity of the Lagrangian (21a), namely we must satisfy
@@ III-B Illustrative Example
We now revisit example (8). In Fig. 2 we display the network given in (8) (Left) and the corresponding solution to (13) (Right). We see that the influence from the measured node 1 to the unobserved node 4 is removed, moreover, since all edges ($A_{4j}$) into node 4 are structurally decoupled they are entirely removed. However, since nodes 2 and 3 influence the dynamics of node 1, we see that structurally Fig. 2 (Right) essentially permutes the role of nodes 2 and 3 when compared to Fig. 2 (Left). We note here that the structurally essential edges $A_{1j}$ remain present, as expected, in both networks. Moreover, the condition in Theorem 2 is satisfied and hence it suffices to solve (18) to find the most dissimilar network. This example illustrate a few key points, namely, (i) nodes that do not influence the measured nodes can be removed from the network, (ii) the relationship between unmeasured but influencing nodes can be changed by a simple permutation of their labels.
Figure 2: Network described by (8) (Left) and the equivalent maximally dissimilar network (solution to (13)) (Right). Blue dashed circles indicate which nodes are measured.
@@ IV Networks with $\epsilon$-Close measurements
In practice, measurements are associated with some uncertainty due to noise, furthermore, executing the same experiment twice will yield similar yet non-identical measurements. We now begin to characterize how measurements that are “close” together influence the set of possible network topologies that are consistent with them. Moving forward, we assume both $A$ and $A+\Delta$ are Hurwitz so that the corresponding observability Gramians are well defined.
To analyze this setting we consider an augmented system akin to [15, 16]. Specifically, we define
$\bar{A}\coloneq\text{blkdiag}(A,A+\Delta),~~\bar{C}\coloneq\begin{bmatrix}C&-C\end{bmatrix}$, (26)
@@
$\bar{\Sigma}:\begin{cases}\dot{\bar{x}}(t)=\bar{A}\bar{x}(t)&\bar{x}(0)=\begin{bmatrix}x_{0}&x_{0}\end{bmatrix}^{\top}\\ e(t)=\bar{C}\bar{x}(t).\end{cases}$ (27)
Notice, $\|e(t)\|_{2}$ is a measure of the cumulative error between measurements of the original and perturbed systems.
@@
Let $\epsilon>0$ and $\bar{W}_{o}$ be the observability Gramian for the augmented system $\bar{\Sigma}$ given in (27), if $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$, then $\|e(t)\|_{2}<\epsilon$.
@@
Assume $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$. Notice
@@
For future use, we establish the following definition.
@@
Analogous to Section III we now consider a similar program to (13) by modifying the constraints. The following program generates the maximally dissimilar network
@@
By optimizing over $\bar{W}_{o}$, this program is nonconvex (bilinear constraint), however, this program can be solved with branch and bound algorithms in YALMIP [17].
Theorem 3 elucidates the relationship between the observability Gramian, $\bar{W}_{o}$, of $\bar{\Sigma}$ (27) and how “close” the measurements of the two systems will be. With this in mind, we can now consider how the set of similar networks is influenced by the spectral properties of $\bar{W}_{o}$.
IV-A Interpretations Based on Observability Gramian
We now consider the implications of (29), which is also a constraint in the optimization program (30). This equation characterizes all possible Gramians $\bar{W}_{o}$ that correspond to networks with $\epsilon$-close measurements. However, given a feasible $\bar{W}_{o}$, there can exist multiple values of $\bar{A}$, which translate to multiple candidate networks via the perturbation $\Delta$. We are interested in further analyzing this relationship. We leverage and extend techniques from [18] to determine what perturbations $\Delta$ are consistent with a given $\bar{W}_{o}$. Specifically, we will explicitly show the dependence of $\Delta$ on the spectra of $\bar{W}_{o}$.
Since $\bar{W}_{o}$ is a symmetric matrix we can express $\bar{W}_{o}=VDV^{\top}$. Denoting $\text{dim}(\mathcal{N}(\bar{W}_{o}))=l$ we have $V=\begin{bmatrix}V_{o}&V_{\bar{o}}\end{bmatrix}$ is a orthogonal matrix with $V_{o}\in\mathbb{R}^{2n\times(2n-l)}$ and $V_{\bar{o}}\in\mathbb{R}^{2n\times l}$ encoding a basis for the observable subspace and the unobservable subspace of $\bar{\Sigma}$ respectively. Here, $D=\text{blkdiag}(\Lambda,0)$ is a diagonal matrix where $\Lambda\in\mathbb{R}^{(2n-l)\times(2n-l)}$ is full rank. Rewriting (29) by pre and post multiplying by $V^{\top}$ and $V$ respectively we arrive at
@@ IV-A Interpretations Based on Observability Gramian
will correspond to an observable decomposition of $\bar{\Sigma}$, (31) succinctly reduces to
@@
Following [18] we can by pre and post multiply by $\Lambda^{-1/2}$ to yield
@@
This implies the symmetric part of $\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}$ is $\frac{1}{2}\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}$. In general we have that
@@
for an arbitrary skew-symmetric matrix $S$. Finally yielding
@@
Now we have that
$\bar{A}=V\left(\begin{bmatrix}-\frac{1}{2}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}+\Lambda^{-1/2}S\Lambda^{1/2}&0\\ \check{A}_{21}&\check{A}_{\bar{o}}\end{bmatrix}\right)V^{\top}.$ (37)
@@
Thus the potentially many solutions, $(S,\check{A}_{21},\check{A}_{\bar{o}})$, to (38) will characterize the set of possible perturbations, $\Delta$, and hence, resulting networks, $A+\Delta$, that generate similar measurements. Moreover, if two inference methods result in two different network hypotheses $A$ and $\tilde{A}$ with different topologies, one could compute the augmented Gramian (i.e., solve (29) with $)\bar{A}=\text{blkdiag}(A,\tilde{A}$)) and solve (38) to find other topologies that fall within the error defined by $A$ and $\tilde{A}$.
We also see from (38) that for a given feasible observability Gramian all networks share a component related to the observable subsystem of $\bar{\Sigma}$ given in (27), and differ by the skew-symmetric matrix $S$ and the two free matrices $\check{A}_{21}$ and $\check{A}_{\bar{o}}$. To illustrate some of the quantitative trends we consider two cases. If $\Lambda_{ii}\ll 1$, the fixed term will be dominant, thus
@@
This implies that only the observable term will contribute to the set of possible perturbed networks, which is consistent with our previous intuition from Section II that the observable portions of the networks must be consistent. Similarly we see that if we tolerate significant deviations in measurements, $\Lambda_{ii}\gg 1$ we have that
$\bar{A}\approx V_{o}\Lambda^{-1/2}S\Lambda^{1/2}V_{o}^{\top}+V_{\bar{o}}\check{A}_{12}V_{o}^{\top}+V_{\bar{o}}\check{A}_{\bar{o}}V_{\bar{o}}^{\top}.$ (40)
Implying that there can be significant variability in the network structure due to the free variables influence.
@@ V-A Identical Measurements with Random Networks
To explore the impacts of partial measurement on topology identification we study edge misclassification in Erdős-Rényi and Watts-Strogatz random networks. The choice to study Watts-Strogatz networks arises from their small-world characteristics that are ubiquitous in brain networks [19]. Specifically, we consider networks with $n=100$ nodes with edge presence probability $p=1/6$, and edge flip probability $\beta=1/30$. For the Watts-Strogatz networks we connect each node to $K=3$ of its neighbors. We then sample from a standard normal distribution to select the edge weights. We randomly select which nodes to measure and compute the corresponding maximally dissimilar networks by solving (16). In each case, we compute the percentage of edges that have been flipped (removed/added), we deemed an edge present if the weight was greater $10^{-5}$. We average the results over $100$ trials with different random seeds, the results are indicated in Fig. 3.
Notably, we observe that when measuring under $6$ nodes the random networks are essentially completely unobservable and the most dissimilar networks effectively change the entire network topologies without affecting the measurements. However, after observing greater than $6$ nodes there is a significant decline in the percentage of edges flipped for both random network models. This indicates a phase transition from completely unobservable networks to completely observable networks.
@@ V-B $\epsilon$-Close Measurements
To explore the implications of the program in (30), we considered the following network and observation scheme
@@
Moreover, to avoid solving the nonconvex problem (30) we simply fixed $\bar{W}_{o}$ to ensure it was feasible, and removed the constraint $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|_{2}^{2}}I$. In particular, we have that $\sqrt{\lambda_{\max}(\bar{W}_{o})}=0.6487$ and the corresponding maximally dissimilar network was
@@
In Fig. 4 we display the observed trajectories, $y(t)$, for the original system, and the most dissimilar system (i.e., solution to (30)), $\tilde{y}(t)$. We selected the initial condition so that $\|x_{0}\|_{2}=1$. As displayed the trajectories remain close over the entire simulation and the error norm was $\|e(t)\|_{2}=0.0357<0.6487\sqrt{2}$. In Fig. 5 we display the corresponding network topologies. We see that the general intuition developed in Section II is violated as node 2, despite influencing the measured node 1, has all connections removed, and node 3 is now compensating by influencing node 1. In Fig. 5 a direct comparison of the most dissimilar networks that produce identical (Right) and close (Middle) is displayed. Allowing the measurements to be “close” yields significant topological variations.
@@
Figure 5: (Left) Original network, (Middle) network that generates similar measurements that is the most dissimilar (solution to (30)), (Right) most dissimilar network that produces identical measurements (solution to (13)). Blue dashed circles indicate which nodes are measured.

### v2

@@ III-A Equivalence of the $\ell_{1}$ and $\ell_{2}$ Programs
Consider the program given in (14); we now relax it to instead minimize the $\ell_{2}$ norm:
@@
This problem is convex; to find the optimal solution, take the gradient of the objective and set it to zero:
@@
Note that we can vectorize this to yield a linear equation, i.e., $(I_{n}\otimes\Phi^{\top})(a+\Gamma v^{*})=0$. Given the solution $V^{*}$ to matrix equation (18), we now provide a condition under which it is also a solution to (13).
@@
Let $V^{*}$ be a minimizer for (17). It is also a minimizer to the original program (13) if
@@
Let $v^{*}=\text{vec}(V^{*})$ where $V^{*}$ is a minimizer of (17). Assume that $V^{*}$ satsifies (19). Recall that the programs (13) and (16) are equivalent. We now show that for an appropriate choice $(v^{*},t^{*},\lambda^{*})$, where $\lambda^{*}=\begin{bmatrix}{\lambda_{1}^{*}}^{\top}&{\lambda_{2}^{*}}^{\top}\end{bmatrix}^{\top}$ are the dual variables associated with the constraints in (16), the Karush-Kuhn-Tucker (KKT) conditions for (16) are satisfied. We select $t^{*}=|a+\Gamma v^{*}|$, and $\lambda_{1,i}^{*}=1$ and $\lambda_{2,i}^{*}=0$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=1$, $\lambda_{1,i}^{*}=0$ and $\lambda_{2,i}^{*}=1$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=-1$, and lastly $\lambda_{1,i}^{*}=1/2=\lambda_{2,i}^{*}$ if $\text{sgn}(a+\Gamma v^{*})]_{i}=0$. With these choices $\lambda_{1}^{*}-\lambda_{2}^{*}=\text{sgn}(a+\Gamma v^{*})$. The Lagrangian of (16) is
@@
$\begin{bmatrix}0\\ \mathds{1}\end{bmatrix}+\begin{bmatrix}\Gamma^{\top}&-\Gamma^{\top}\\ -I_{n^{2}}&-I_{n^{2}}\end{bmatrix}\lambda^{*}=0$ (21a)
@@
With $t^{*}$ as chosen, the feasibility condition (21b) is satisfied. Moreover, (21c) is satisfied by our choice of $\lambda^{*}$. It remains to check (21a) and (21d). Notice, from the stationarity of the Lagrangian (21a) we have that $\lambda_{1}^{*}+\lambda_{2}^{*}=\mathds{1}$ which is also satisfied from our choice of $\lambda^{*}$. Rewriting (21d) as
@@
Since $\lambda_{1}^{*}-\lambda_{2}^{*}=\text{sgn}(a+\Gamma v^{*})$ these equations are satisfied, and thus (21d) is satisfied. Lastly, we must consider the first n(nr)n(n-r) rows of the stationarity of the Lagrangian (21a),
@@ III-B Illustrative Example
We now revisit example (8). In the left half of Fig. 2, we display the original network (i.e., (8)); in the right half of the figure, we display the minimizer to (13), which is the most structurally dissimilar network that still exhibits identical measurements as the original network. In this example, the condition in Theorem 2 is satisfied, so it suffices to solve (18) to find the most dissimilar network.
Not counting self-dynamics, the two networks have only two edges in common. Since all edges ($A_{4j}$) into node 4 are structurally decoupled, they are entirely removed in the dissimilar network. Notice that since nodes 2 and 3 influence the dynamics of node 1, the dissimilar network essentially permutes the role of nodes 2 and 3 compared to the original network. Conversely, structurally essential edges $A_{1j}$ remain present in both networks, as expected.
This example illustrate a few key points. First, edges connected to nodes that do not influence measured nodes can be removed from the network. Second, the relationship between unmeasured but influential nodes can be changed by a simple permutation of their labels.
Figure 2: (Left) Network described by (8). (Right) The most structurally dissimilar network exhibiting identical measurements as the original network. Blue dashed circles indicate measured nodes.
@@ IV Networks with $\epsilon$-Close Measurements
In practice, measurements are corrupted by noise, and executing the same experiment twice will yield similar yet non-identical measurements. We now extend previous analysis to account for this: we characterize the set of networks who yield similar (not necessarily identical) measurements to the original system. We assume both $A$ and $A+\Delta$ are Hurwitz [^2] In practice this assumption is necessary as unstable systems would yield exponentially growing trajectories; it is less meaningful to compare the “closeness” of such trajectories. so that the corresponding observability Gramians are well-defined. We consider an augmented system akin to [16, 1]:
$\bar{A}=\text{blkdiag}(A,A+\Delta),~~\bar{C}=\begin{bmatrix}C&-C\end{bmatrix}$, (26)
@@
$\bar{\Sigma}:\begin{cases}\dot{\bar{x}}(t)=\bar{A}\bar{x}(t),&\bar{x}(0)=\begin{bmatrix}x_{0}^{\top}&x_{0}^{\top}\end{bmatrix}^{\top}\\ e(t)=\bar{C}\bar{x}(t).\end{cases}$ (27)
Here, $\|e(t)\|_{2}$ is a measure of the cumulative error between measurements of the original and perturbed systems.
@@
Let $\epsilon>0$ and $\bar{W}_{o}$ be the observability Gramian for the augmented system $\bar{\Sigma}$ given in (27).
If $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$, then $\|e(t)\|_{2}<\epsilon$.
@@
Assume $\bar{W}_{o}\prec\frac{\epsilon^{2}}{2\|x_{0}\|^{2}_{2}}I$. Then,
@@
Theorem 3 elucidates the relationship between $\bar{W}_{o}$, the observability Gramian of (27), and how close the measurements of the two systems will be. Now, we leverage it to consider the most structurally dissimilar network that produces $\epsilon$-close measurements as the original network. This is similar to Section III; however, we no longer require identical measurements. We first establish the following definition.
@@
The following program generates the maximally dissimilar network (analogous to (13)) for $\epsilon$-close measurements:
@@
This program is nonconvex due to its bilinear constraint, but can be solved with branch and bound algorithms in YALMIP [13].
We now consider the implications of (29). This equation characterizes all possible Gramians $\bar{W}_{o}$ that correspond to networks with $\epsilon$-close measurements to the original. However, given a feasible $\bar{W}_{o}$ there can exist multiple values of $\bar{A}$, which translate to multiple candidate networks via the perturbation $\Delta$. We extend techniques from [8] to determine what perturbations $\Delta$ are consistent with a given $\bar{W}_{o}$.
Since $\bar{W}_{o}$ is a real symmetric matrix, we can diagonalize it using an orthonormal matrix $V$: $\bar{W}_{o}=VDV^{\top}$. We can partition $V$ as $V=\begin{bmatrix}V_{o}&V_{\bar{o}}\end{bmatrix}$, where $l:=\text{dim}(\mathcal{N}(\bar{W}_{o}))$, $V_{o}\in\mathbb{R}^{2n\times(2n-l)}$, and $V_{\bar{o}}\in\mathbb{R}^{2n\times l}$. The columns of $V_{o}$ and $V_{\bar{o}}$ form a basis for the observable and the unobservable subspace of $\bar{\Sigma}$, respectively. Diagonal matrix $D$ can be partitioned as $D=\text{blkdiag}(\Lambda,0)$ where $\Lambda\in\mathbb{R}^{(2n-l)\times(2n-l)}$ is full rank. We now pre- and post-multiply (29) by $V^{\top}$ and $V$:
@@
correspond to an observability decomposition of $\bar{\Sigma}$; so (31) reduces to
@@
Following [8], pre- and post-multiply by $\Lambda^{-1/2}$ to obtain
@@
This implies that the symmetric part of $\Lambda^{1/2}\check{A}_{o}\Lambda^{-1/2}$ is $\frac{1}{2}\Lambda^{-1/2}\check{C}_{o}^{\top}\check{C}_{o}\Lambda^{-1/2}$. In general, we have that
@@
for an arbitrary skew-symmetric matrix $S$. Pre- and post-multiply again to yield
@@
Returning to the full augmented system, we have
$\bar{A}=V\left(\begin{bmatrix}-\frac{1}{2}\Lambda^{-1}\check{C}_{o}^{\top}\check{C}_{o}+\Lambda^{-1/2}S\Lambda^{1/2}&0\\ \check{A}_{21}&\check{A}_{\bar{o}}\end{bmatrix}\right)V^{\top},$ (37)
@@
From this analysis, we see that solutions $(S,\check{A}_{21},\check{A}_{\bar{o}})$ to (38) will characterize perturbations $\Delta$ and resulting networks $A+\Delta$ that generate similar measurements.
We also see from (38) that for a given feasible observability Gramian, all networks share a component related to the observable subsystem of $\bar{\Sigma}$ given in (27), and differ by the skew-symmetric matrix $S$ and the two free matrices $\check{A}_{21}$ and $\check{A}_{\bar{o}}$. To illustrate some of the quantitative trends, we consider two cases. If $\Lambda_{ii}\ll 1$, the fixed term will be dominant, thus
@@
This implies that only the observable term will contribute to the set of possible perturbed networks, which is consistent with our previous intuition from Section II that the observable portions of the networks must be consistent. Similarly we see that if we tolerate significant deviations in measurements, $\Lambda_{ii}\gg 1$ we have
$\bar{A}\approx V_{o}\Lambda^{-1/2}S\Lambda^{1/2}V_{o}^{\top}+V_{\bar{o}}\check{A}_{12}V_{o}^{\top}+V_{\bar{o}}\check{A}_{\bar{o}}V_{\bar{o}}^{\top},$ (40)
which implies that there can be significant variability in the network structure due to the free variables’ influence.
@@ V-A Identical Measurements with Random Networks
To explore the impact of partial measurement on network structure identification, we study edge misclassification in Erdős-Rényi and Watts-Strogatz random networks; the latter exhibits properties similar to brain networks [2]. We consider networks with $n=100$ nodes with edge presence probability $p=1/6$ and edge flip probability $\beta=1/30$. For the Watts-Strogatz networks, we connect each node to $K=3$ of its neighbors. We sample from a standard normal distribution to select edge weights, and randomly select which nodes to measure.
We compute the maximally dissimilar networks (that exhibit identical dynamics) by solving (16). In each case, we compute the percentage of edges that have been flipped (removed/added) relative to the original network; we deemed an edge present if the weight was greater $10^{-5}$. Results were averaged over $100$ trials with different random seeds and are depicted in Fig. 3. When measuring under $6$ nodes, the most dissimilar networks effectively change the entire network structure without affecting the measurements. However, after observing greater than $6$ nodes, there is a significant decline in the percentage of edges flipped for both network models. This indicates a phase transition from completely “unobservable” networks to completely “observable” networks.
@@ V-B $\epsilon$-Close Measurements
To explore the implications of the program in (30), we consider the following network and observation scheme
@@
For example’s sake, we seek an exact solution; thus, instead of solving nonconvex problem (30), we work backwards. We fix $\bar{W}_{o}$ to ensure feasibility and temporarily remove the other constraint. Then, we find that $\sqrt{\lambda_{\max}(\bar{W}_{o})}=0.65$ and the corresponding maximally structurally dissimilar network is
@@
In Fig. 4, we show the observed trajectories $y(t)$ and $\tilde{y}(t)$ for the original system and most dissimilar system, respectively. We selected the initial condition so that $\|x_{0}\|_{2}=1$. We see that the trajectories remain close over the entire simulation; the error norm is $\|e(t)\|_{2}=0.036$. This is less than $0.65\sqrt{2}$, which is the upper bound provided by Theorem 3. In Fig. 5, we display the corresponding network structures. We see that the general intuition developed in Section II is violated as node 2, despite influencing measured node 1, has all connections removed; this is compensated by node 3’s influence on node 1. For comparison, we also include the most dissimilar network that produces identical measurements; we see that allowing the measurements to be close instead of identical yields significant structural variations.
@@
Figure 5: (Left) Original network. (Middle) The most structurally dissimilar network that generates similar measurements. (Right) The most structurally dissimilar network that generates identical measurements. Blue dashed circles indicate which nodes are measured.